OpenStudy (anonymous):

solve; find all real and imaginary solutions in exact form. x^4-9x^2-10=0

OpenStudy (anonymous):

maybe easier to start with \[u^2-9u-10=0\] find the two solutions, then replace \(u\) by \(x^2\) and solve for \(x\)

OpenStudy (anonymous):

find the dicriminent then the roots of 4x^2-2x+4=0

OpenStudy (anonymous):

not to bad since \[u^2-9u-10\] factors as \[(u-10)(u+1)\] so you can easily solve \[(u-10)(u+1)=0\]

OpenStudy (anonymous):

that makes sense!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so I got 10 and -1 but when you plug both of them back in, neither of them work

OpenStudy (anonymous):

of course they do they work for \[u^2-9u-10=0\] but not for the original equation now you have to solve \[x^2=10\] and also \(x^2=-1\)

OpenStudy (anonymous):

ohhhh! I see ok

OpenStudy (anonymous):

both are easy, the second one gives you the two complex solutions

OpenStudy (anonymous):

yea, I see what I did wrong, I plugged it into the wrong equation. thanks!

OpenStudy (anonymous):

so to find \[x^2=10\] and \[x^2=-1\] you have to take the square root of each side?